OpenStudy (anonymous):

The measure of a side of a square field is . What is its perimeter? (Points : 1)

OpenStudy (angela210793):

P=4*side=4*1=4

OpenStudy (katrinakaif):

4(n)

OpenStudy (anonymous):

4*1=4

OpenStudy (anonymous):

Thanks Katrina

OpenStudy (katrinakaif):

=D Anytime!

OpenStudy (anonymous):

2x^2 + 3

OpenStudy (anonymous):

that's what the square fieild is

OpenStudy (angela210793):

i didn't get it O.o